# The Symmetry Taco: Equivalences between Gapped, Gapless, and Mixed-State SPTs

### v1

@@ IV From Gapless SPTs to Gapped SPTs via the SymTaco
In this Section, we will leverage the SymTFT framework reviewed in the prior two Sections to establish our first result: namely, a correspondence between $1+1$d $G$-symmetric gapless SPTs and certain $G\times G$-symmetric gapped phases. This correspondence follows from the SymTaco  construction, whereby the SymTFTs of the aforementioned families of gapless and gapped phases can be directly related via the folding and cutting procedures displayed in Fig. 1. Here, the folded bulk $\mathcal{D}(G)$ TO, which is folded along a subregion which hosts $\mathcal{D}(G)/\mathcal{A}$ TO, defines the symmetry taco alluded to in Section I. While our primary objective in introducing the SymTaco  is for extending the symTFT perspective to mixed-state phases (which we develop in the following Section V), it also sheds new light on the emergence of gapped $G\times G$-symmetric phases in certain tensor network descriptions of $\mathcal{D}(G)$ TOs [81] and provides a systematic approach for relating gapless and gapped SPTs.
@@
Let us first construct the SymTaco  for general anyon theories: we begin with a $2+1$d TO whose anyon theory is given by a unitary modular tensor category $\mathcal{C}$ and condense anyons belonging to a condensable subalgebra $\mathcal{A}$ in some subregion. This reduces the TO in the subregion to $\mathcal{C}/\mathcal{A}$, which is separated from the $\mathcal{C}$ TO by a pair of gapped domain walls, which we assume are identical, i.e., they are described by the same condensable subalgebras. This choice endows the system with reflection symmetry about an appropriately chosen mirror line passing through the condensed subregion, which maps the two regions with $\mathcal{C}$ TO onto one another. We obtain the SymTaco  by folding the system along this mirror line, yielding a bilayer with TO given by $\mathcal{C}\boxtimes\overline{\mathcal{{C}}}$ [^8] Here, $\overline{\mathcal{C}}$ denotes the unitary modular tensor category with the reversed braiding of $\mathcal{C}$, i.e., it is the time-reversed conjugate of $\mathcal{C}$. which has a gapped boundary to vacuum along the fold. More precisely, starting from the doubled bulk TO and moving towards the fold, there is a gapped boundary from $\mathcal{C}\boxtimes\overline{\mathcal{{C}}}$ to $\left(\mathcal{C}\boxtimes\overline{\mathcal{{C}}}\right)/\left(\mathcal{A}\boxtimes\bar{\mathcal{A}}\right)$, followed by a gapped boundary from the latter TO to vacuum. It is this particular setup that we refer to as the SymTaco.
@@
The key point, which we will expound on shortly in this Section, is that we can only obtain a subset of all possible gapped boundaries between the doubled TO $\mathcal{C}\boxtimes\overline{\mathcal{{C}}}$ and vacuum when the former is obtained via the folding procedure described above. We will refer to this restricted set of gapped boundaries as folded boundaries. As reviewed in Sec. II, in the case when $\mathcal{C}=\mathcal{D}(G)$ these gapped boundaries correspond to gapped phases of $1+1$d $G\times G$-symmetric systems; hence, we expect the condensable subalgebras $\mathcal{A}$ to correspond to a subset of all $G\times G$-symmetric gapped phases in $1+1$d. On the other hand, we can also imagine breaking the SymTaco  in half (or cutting the unfolded theory along the mirror line), in which case the resulting edge would support a gapless $G$-SPT whose symmetry data is equivalently encoded in the condensable subalgebra $\mathcal{A}$, as discussed in Sec. III. Through the condensable subalgebra $\mathcal{A}$, we can thus obtain a one-to-one correspondence between $G$-igSPTs and a subset of gapped $G\times G$-symmetric $1+1$d phases. In the following subsections, we work out the explicit constraints on the condensable subalgebras that characterize folded boundaries of general bilayer topological orders, and then apply these constraints to the case of $\mathcal{D}(G)\boxtimes\overline{\mathcal{D}(G)}\cong\mathcal{D}(G\times G)$. This will allow us to work out the subgroup $H\subset G\times G$ and 2-cocycle $\omega\in H^{2}(H,U(1))$ characterizing such boundaries and to show that they do in fact correspond to the data defining a $G$-igSPT.
@@ IV.1 Characterization of Folded Boundaries
Here, we discuss folded boundaries of general $2+1$d TOs given by $\mathcal{C}\boxtimes\overline{\mathcal{{C}}}$ and specialize to the case of interest $\mathcal{C}=\mathcal{D}(G)$ in the following subsection. For a general gapped boundary to vacuum of the $\mathcal{C}\boxtimes\overline{\mathcal{{C}}}$ TO, we will show that whether or not it corresponds to a folded boundary of the SymTaco  is determined by the positivity of a certain matrix, which encodes the data of the condensed anyons in the condensable subalgebra $\mathcal{A}$. As it turns out, this positivity constraint is closely related to the positivity constraint on physical density matrices, which we leverage in Section V to obtain the SymTaco  construction for mixed-state SPTs in $1+1$d.
The data for a general $2+1$d TO is specified by a unitary modular tensor category $\mathcal{D}$, which contains the (finite) set of topologically distinct anyons (or superselection sectors) of the theory, $\{a\}$. We denote the vacuum (or trivial) superselection sector as $1$ and denote the inverse of an anyon $a\in\mathcal{D}$ as $\bar{a}$, where $\bar{a}$ is the unique anyon in $\mathcal{D}$ such that its fusion with $a$ contains the vacuum as a fusion product: $a\times\bar{a}=1+\dots$.
@@
Let us turn to the object of interest: a bilayer Abelian TO of the form $\mathcal{D}=\mathcal{C}\boxtimes\bar{\mathcal{C}}$. In this case, the anyons of $\mathcal{D}$ can be written as ordered pairs $(a,b)$ (for $a,b\in\mathcal{C}$, with the understanding that anyons in $\bar{\mathcal{C}}$ have opposite braiding to those in $\mathcal{C}$). In general, the distinct gapped boundaries of this TO are classified via the data provided in Sec. II.2. Suppose now though that we are considering a SymTaco, i.e., we started with the TO given by $\mathcal{C}$, condensed anyons in accordance with some condensable subalgebra $\mathcal{A}$ in a middle region, and then performed the fold along the mirror line to obtain the folded boundary of the SymTaco. In this case, the indicator function $M$ corresponding to the folded boundary obeys the following constraints:
$M_{(a,b)}=1\implies\begin{array}[]{ll}&M_{(b,a)}=1\,,\\ &M_{(a\times b,1)}=1.\end{array}$ (36)
@@
The first constraint arises from the layer exchange symmetry of the SymTaco, which descends from the reflection symmetry about the fold line in the unfolded theory. The second constraint can likewise be understood from the unfolded picture: if $M_{(a,b)}=1$, in the unfolded theory this implies that if we bring an anyon $a$ from the left region (with $\mathcal{C}$ TO) and an anyon $b$ from the right region (with $\mathcal{C}$ TO) towards the fold line, either (i) both $a,b$ condense at the gapped boundaries between $\mathcal{C}$ and $\mathcal{C}/\mathcal{A}$, or (ii) they both pass into the region $\mathcal{C}/\mathcal{A}$ and annihilate each other. Suppose now that we instead bring both $a$ and $b$ from the left region with $\mathcal{C}$ TO and bring them towards the middle region with $\mathcal{C}/\mathcal{A}$ TO. Then, either (i) both condense at the gapped boundary to $\mathcal{C}/\mathcal{A}$ or, (ii) they annihilate each other after passing through this gapped boundary. In the folded theory, this implies that $M_{(a\times b,1)}=1$. Note that, by combining Eq. (36) with Eq. (35), we can also derive from $M_{(a,b)}=1$ that $M_{(1,a\times b)}=1$ and $M_{(a,\bar{a})}=M_{(b,\bar{b})}=1$.
@@
In fact, the constraints in Eq. (36) provide a sufficient condition for a general gapped boundary of $\mathcal{D}$ to also define a consistent folded boundary. In other words, given a $\mathcal{C}\boxtimes\bar{\mathcal{C}}$ TO with a gapped boundary whose indicator function $M$ satisfies Eq. (36), we can derive a condensable subalgebra $\mathcal{A}$ of $\mathcal{C}$ such that this boundary can equivalently be realized as a folded boundary of the corresponding SymTaco. Indeed, we can define the set of condensed anyons $\mathcal{A}$ such that an anyon $c\in\mathcal{C}$ is in $\mathcal{A}$ if $M_{(c,1)}=1$. That $\mathcal{A}$ is a condensable subalgebra follows immediately from the fact that $(c,1)$ are elements of a Lagrangian subalgebra, and thus all $c\in\mathcal{A}$ must have trivial topological spin and trivial braiding with all other elements of $\mathcal{A}$. Given this, we can straightforwardly construct a folded boundary by taking a $\mathcal{C}$ TO, picking the condensable subalgebra $\mathcal{A}$ we have constructed above, condensing it in a subregion, and folding along the mirror symmetric line; this procedure automatically realizes a folded boundary of $\mathcal{C}\boxtimes\bar{\mathcal{C}}$ to vacuum. One may check that this folded boundary is precisely the original gapped boundary we started with.
@@
From the above discussion, we see that a gapped boundary of $\mathcal{C}\boxtimes\bar{\mathcal{C}}$ can be realized as a folded boundary if and only if Eq. (36) holds. This constraint can be more compactly expressed as imposing a positivity condition on the indicator function $M$. First, by combining Eqs. (35) and (36), we can derive the following inequality:
$M_{(a,\bar{b})}+M_{(a,\bar{c})}\leq M_{(a,\bar{a})}+M_{(b,\bar{c})}\,.$ (37)
@@
Conversely, if this inequality holds, then so does Eq. (36) (see Appendix B for the proof). We now make use of the following fact: a symmetric binary matrix $M$ is positive semi-definite if and only if Eq. (37) holds [95]. Interpreting the indicator function $M_{(a,\bar{b})}$ as such a symmetric binary matrix, it follows from Eq. (37) that the indicator function $M_{(a,\bar{b})}$ is positive semi-definite. Therefore, a gapped boundary of $\mathcal{C}\boxtimes\bar{\mathcal{C}}$ can be realized as a folded boundary of the SymTaco  if and only if the corresponding indicator function $M$ is a symmetric, positive semi-definite matrix [^9] Positivity also played an important role in the derivation by Ref. [81], which established a similar correspondence between folded boundaries of quantum doubles and $G\times G$-SPTs..
@@
To illustrate the constraint that the requirement of being a folded boundary imposes on general gapped boundaries, consider the boundary of $\mathcal{D}(\mathbb{Z}_{2})\times\mathcal{D}(\mathbb{Z}_{2})\cong\mathcal{D}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$ that corresponds to the nontrivial $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ $1+1$d SPT. The anyons condensed at the boundary are $1$, $e_{1}m_{2}$, $m_{1}e_{2}$, and $f_{1}f_{2}$ (see Section II.2). We can immediately see that the condition of Eq. (37) is not satisfied: we have $M_{(e,m)}=M_{(m,e)}=1$ but $M_{(e\times m,1)}=0$. Therefore, the positivity condition is violated and this boundary cannot be obtained via folding, i.e., while this is valid gapped boundary to vacuum of a general $\mathcal{D}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$ TO, it cannot describe a gapped boundary to vacuum for a $\mathcal{D}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$ TO that is obtained by folding a $\mathcal{D}(\mathbb{Z}_{2})$ TO along a subregion where some condensable subalgebra $\mathcal{A}$ is condensed.
@@
such that $M_{a}$ is the multiplicity of fusion channels of the anyon $a$ when fused to the boundary. The anyon $a$ is confined (i.e. not condensed) at the boundary if $M_{a}=0$. This reduces to the definition provided above in the case of Abelian TOs–Abelian anyons can only have multiplicity 00 or $1$ [96].
@@
We now apply the folding trick to obtain the SymTaco, as shown in Fig. 1, which results in a gapped boundary to vacuum for the $\mathcal{C}\boxtimes\bar{\mathcal{C}}$ TO. Importantly, anyons from this TO undergo a sequence of two condensations to get to the boundary: first, an anyon $(a,b)\in\mathcal{C}\boxtimes\bar{\mathcal{C}}$ tunnels into the region with $\mathcal{C}^{\prime}\boxtimes\bar{\mathcal{C}}^{\prime}$ TO and turns into $\bigoplus_{\alpha\beta}W_{a\alpha}W_{b\beta}\,(\alpha,\beta)$. Next, anyons from $\mathcal{C}^{\prime}\boxtimes\bar{\mathcal{C}}^{\prime}$ can condense at the gapped boundary to vacuum specified by condensation of the Lagrangian subalgebra [^10] Following Ref. [98], we use the same labels for objects in $\mathcal{C^{\prime}}$ and $\mkern 1.5mu\overline{\mkern-1.5mu\mathcal{C}\mkern-1.5mu}\mkern 1.5mu^{\prime}$, such that $\mkern 1.5mu\overline{\mkern-1.5mu\alpha\mkern-1.5mu}\mkern 1.5mu$ denotes the dual of $\alpha.$
$\mathcal{L}^{\prime}=\bigoplus_{\alpha}\alpha\bar{\alpha}\,,$ (40)
which corresponds to the canonical gapped boundary of the condensed TO $\mathcal{C}^{\prime}\boxtimes\overline{\mathcal{C}}^{\prime}$. In other words, the corresponding condensation matrix for this gapped boundary to vacuum is given by $M^{\prime}_{(\alpha,\bar{\beta})}=\delta_{\alpha\beta}$. Hence, the gapped boundary to vacuum of the SymTaco, i.e., of the $\mathcal{C}\boxtimes\bar{\mathcal{C}}$ TO is specified by the data describing this sequence of condensations. The condensation matrix $M$ describing the folded boundary of the SymTaco  is therefore given by
$M_{(a,\bar{b})}=W_{a,\alpha}W_{\bar{b},\bar{\beta}}\delta_{\alpha\bar{\beta}}=W_{a,\alpha}W_{\bar{b},\alpha}=(WW^{T})_{a,\bar{b}}\,,$ (41)
from which we see that $M_{(a,\bar{b})}$ is a positive semidefinite matrix. Thus, the general condition for when a gapped boundary of an arbitrary TO $\mathcal{C}$ provides an admissible folded boundary of the corresponding SymTaco  is encoded in the positivity of the condensation matrix $M$, which is constructed through the intermediate step of the tunneling matrix $W$ (which is in turn specified by the condensable subalgebra $\mathcal{A}$ that is part of the data of the SymTaco  as defined in this paper).

### v2

@@ IV From Gapless SPTs to Gapped SPTs via the SymTaco
In this Section, we will leverage the SymTFT framework reviewed in the prior two Sections to establish our first result: namely, a correspondence between 1+11+1d $G$-symmetric gapless SPTs and certain $G\times G$-symmetric gapped phases. This correspondence follows from the SymTaco  construction, whereby the SymTFTs of the aforementioned families of gapless and gapped phases can be directly related via the folding and cutting procedures displayed in Fig. 1. Here, the folded bulk $\mathcal{D}(G)$ TO, which is folded along a subregion which hosts $\mathcal{D}(G)/\mathcal{A}$ TO, defines the symmetry taco alluded to in Section I. While our primary objective in introducing the SymTaco  is for extending the symTFT perspective to mixed-state phases (which we develop in the following Section V), it also sheds new light on the emergence of gapped $G\times G$-symmetric phases in certain tensor network descriptions of $\mathcal{D}(G)$ TOs [86] and provides a systematic approach for relating gapless and gapped SPTs.
@@
Let us first construct the SymTaco  for general anyon theories: we begin with a 2+12+1d TO whose anyon theory is given by a unitary modular tensor category $\mathcal{C}$ and condense anyons belonging to a condensable subalgebra $\mathcal{A}$ in some subregion. This reduces the TO in the subregion to $\mathcal{C}/\mathcal{A}$, which is separated from the $\mathcal{C}$ TO by a pair of gapped domain walls, which we assume are identical, i.e., they are described by the same condensable subalgebras. This choice endows the system with reflection symmetry about an appropriately chosen mirror line passing through the condensed subregion, which maps the two regions with $\mathcal{C}$ TO onto one another. We obtain the SymTaco  by folding the system along this mirror line, yielding a bilayer with TO given by $\mathcal{C}\boxtimes\overline{\mathcal{{C}}}$ [^9] Here, $\overline{\mathcal{C}}$ denotes the unitary modular tensor category with the reversed braiding of $\mathcal{C}$, i.e., it is the time-reversed conjugate of $\mathcal{C}$. which has a gapped boundary to vacuum along the fold. More precisely, starting from the doubled bulk TO and moving towards the fold, there is a gapped boundary from $\mathcal{C}\boxtimes\overline{\mathcal{{C}}}$ to $\left(\mathcal{C}\boxtimes\overline{\mathcal{{C}}}\right)/\left(\mathcal{A}\boxtimes\bar{\mathcal{A}}\right)$, followed by a gapped boundary from the latter TO to vacuum. It is this particular setup that we refer to as the SymTaco.
@@
The key point, which we will expound on shortly in this Section, is that we can only obtain a subset of all possible gapped boundaries between the doubled TO $\mathcal{C}\boxtimes\overline{\mathcal{{C}}}$ and vacuum when the former is obtained via the folding procedure described above. We will refer to this restricted set of gapped boundaries as folded boundaries. As reviewed in Sec. II, in the case when $\mathcal{C}=\mathcal{D}(G)$ these gapped boundaries correspond to gapped phases of 1+11+1d $G\times G$-symmetric systems; hence, we expect the condensable subalgebras $\mathcal{A}$ to correspond to a subset of all $G\times G$-symmetric gapped phases in 1+11+1d. On the other hand, we can also imagine breaking the SymTaco  in half (or cutting the unfolded theory along the mirror line), in which case the resulting edge would support a gapless $G$-SPT whose symmetry data is equivalently encoded in the condensable subalgebra $\mathcal{A}$, as discussed in Sec. III. Through the condensable subalgebra $\mathcal{A}$, we can thus obtain a one-to-one correspondence between $G$-igSPTs and a subset of gapped $G\times G$-symmetric 1+11+1d phases. In the following subsections, we work out the explicit constraints on the condensable subalgebras that characterize folded boundaries of general bilayer topological orders, and then apply these constraints to the case of $\mathcal{D}(G)\boxtimes\overline{\mathcal{D}(G)}\cong\mathcal{D}(G\times G)$. This will allow us to work out the subgroup $H\subset G\times G$ and 2-cocycle $\omega\in H^{2}(H,U(1))$ characterizing such boundaries and to show that they do in fact correspond to the data defining a $G$-igSPT.
@@ IV.1 Characterization of Folded Boundaries
Here, we discuss folded boundaries of general 2+12+1d TOs given by $\mathcal{C}\boxtimes\overline{\mathcal{{C}}}$ and specialize to the case of interest $\mathcal{C}=\mathcal{D}(G)$ in the following subsection. For a general gapped boundary to vacuum of the $\mathcal{C}\boxtimes\overline{\mathcal{{C}}}$ TO, we will show that whether or not it corresponds to a folded boundary of the SymTaco  is determined by the positivity of a certain matrix, which encodes the data of the condensed anyons in the condensable subalgebra $\mathcal{A}$. As it turns out, this positivity constraint is closely related to the positivity constraint on physical density matrices, which we leverage in Section V to obtain the SymTaco  construction for mixed-state SPTs in 1+11+1d.
The data for a general 2+12+1d TO is specified by a unitary modular tensor category $\mathcal{D}$, which contains the (finite) set of topologically distinct anyons (or superselection sectors) of the theory, $\{a\}$. We denote the vacuum (or trivial) superselection sector as $1$ and denote the inverse of an anyon $a\in\mathcal{D}$ as $\bar{a}$, where $\bar{a}$ is the unique anyon in $\mathcal{D}$ such that its fusion with $a$ contains the vacuum as a fusion product: $a\times\bar{a}=1+\dots$.
@@
Let us turn to the object of interest: a bilayer Abelian TO of the form $\mathcal{D}=\mathcal{C}\boxtimes\overline{\mathcal{C}}$. In this case, the anyons of $\mathcal{D}$ can be written as $a_{+}b_{-}$ for $a,b\in\mathcal{C}$, with the understanding that anyons with the - label have braiding opposite to their counterparts in $\mathcal{C}$. For notational convenience, in the following we denote the indicator function $M_{a,b}\equiv M_{a_{+}b_{-}}$. In general, the distinct gapped boundaries of this TO are classified via the data provided in Sec. II.2. Suppose now though that we are considering a SymTaco, i.e., we started with the TO given by $\mathcal{C}$, condensed anyons in accordance with some condensable subalgebra $\mathcal{A}$ in a middle region, and then performed the fold along the mirror line to obtain the folded boundary of the SymTaco. In this case, the indicator function $M$ corresponding to the folded boundary obeys the following constraints:
$M_{a,b}=1\implies\begin{array}[]{ll}&M_{b,a}=1\,,\\ &M_{a\times b,1}=1.\end{array}$ (36)
@@
The first constraint arises from the layer exchange symmetry of the SymTaco, which descends from the reflection symmetry about the fold line in the unfolded theory. The second constraint can likewise be understood from the unfolded picture: if $M_{a,b}=1$, in the unfolded theory this implies that if we bring an anyon $a$ from the left region (with $\mathcal{C}$ TO) and an anyon $b$ from the right region (with $\mathcal{C}$ TO) towards the fold line, either (i) both $a,b$ condense at the gapped boundaries between $\mathcal{C}$ and $\mathcal{C}/\mathcal{A}$, or (ii) they both pass into the region $\mathcal{C}/\mathcal{A}$ and annihilate each other. Suppose now that we instead bring both $a$ and $b$ from the left region with $\mathcal{C}$ TO and bring them towards the middle region with $\mathcal{C}/\mathcal{A}$ TO. Then, either (i) both condense at the gapped boundary to $\mathcal{C}/\mathcal{A}$ or, (ii) they annihilate each other after passing through this gapped boundary. In the folded theory, this implies that $M_{a\times b,1}=1$. Note that, by combining Eq. (36) with Eq. (35), we can also derive from $M_{a,b}=1$ that $M_{1,a\times b}=1$ and $M_{a,\bar{a}}=M_{b,\bar{b}}=1$.
@@
In fact, the constraints in Eq. (36) provide a sufficient condition for a general gapped boundary of $\mathcal{D}$ to also define a consistent folded boundary. In other words, given a $\mathcal{C}\boxtimes\overline{\mathcal{C}}$ TO with a gapped boundary whose indicator function $M$ satisfies Eq. (36), we can derive a condensable subalgebra $\mathcal{A}$ of $\mathcal{C}$ such that this boundary can equivalently be realized as a folded boundary of the corresponding SymTaco. Indeed, we can define the set of condensed anyons $\mathcal{A}$ such that an anyon $c\in\mathcal{C}$ is in $\mathcal{A}$ if $M_{c,1}=1$. That $\mathcal{A}$ is a condensable subalgebra follows immediately from the fact that $(c,1)$ are elements of a Lagrangian subalgebra, and thus all $c\in\mathcal{A}$ must have trivial topological spin and trivial braiding with all other elements of $\mathcal{A}$. Given this, we can straightforwardly construct a folded boundary by taking a $\mathcal{C}$ TO, picking the condensable subalgebra $\mathcal{A}$ we have constructed above, condensing it in a subregion, and folding along the mirror symmetric line; this procedure automatically realizes a folded boundary of $\mathcal{C}\boxtimes\overline{\mathcal{C}}$ to vacuum. One may check that this folded boundary is precisely the original gapped boundary we started with.
@@
From the above discussion, we see that a gapped boundary of $\mathcal{C}\boxtimes\overline{\mathcal{C}}$ can be realized as a folded boundary if and only if Eq. (36) holds. This constraint can be more compactly expressed as imposing a positivity condition on the indicator function $M$. First, by combining Eqs. (35) and (36), we can derive the following inequality:
$M_{a,\bar{b}}+M_{a,\bar{c}}\leq M_{a,\bar{a}}+M_{b,\bar{c}}\,.$ (37)
@@
Conversely, if this inequality holds, then so does Eq. (36) (see Appendix C for the proof). We now make use of the following fact: a symmetric binary matrix $M$ is positive semi-definite if and only if Eq. (37) holds [100]. Interpreting the indicator function $M_{a,\bar{b}}$ as such a symmetric binary matrix, it follows from Eq. (37) that the indicator function $M_{a,\bar{b}}$ is positive semi-definite. Therefore, a gapped boundary of $\mathcal{C}\boxtimes\overline{\mathcal{C}}$ can be realized as a folded boundary of the SymTaco  if and only if the corresponding indicator function $M$ is a symmetric, positive semi-definite matrix [^10] Positivity also played an important role in the derivation by Ref. [86], which established a similar correspondence between folded boundaries of quantum doubles and $G\times G$-SPTs..
@@
To illustrate the constraint that the requirement of being a folded boundary imposes on general gapped boundaries, consider the boundary of $\mathcal{D}(\mathbb{Z}_{2})\times\mathcal{D}(\mathbb{Z}_{2})\cong\mathcal{D}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$ that corresponds to the nontrivial $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 1+11+1d SPT. The anyons condensed at the boundary are $1$, $e_{1}m_{2}$, $m_{1}e_{2}$, and $f_{1}f_{2}$ (see Section II.2). We can immediately see that the condition of Eq. (37) is not satisfied: we have $M_{e,m}=M_{m,e}=1$ but $M_{e\times m,1}=0$. Therefore, the positivity condition is violated and this boundary cannot be obtained via folding, i.e., while this is valid gapped boundary to vacuum of a general $\mathcal{D}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$ TO, it cannot describe a gapped boundary to vacuum for a $\mathcal{D}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$ TO that is obtained by folding a $\mathcal{D}(\mathbb{Z}_{2})$ TO along a subregion where some condensable subalgebra $\mathcal{A}$ is condensed.
@@
such that $M_{a}$ is the multiplicity of fusion channels of the anyon $a$ when fused to the boundary. The anyon $a$ is confined (i.e. not condensed) at the boundary if $M_{a}=0$. This reduces to the definition provided above in the case of Abelian TOs–Abelian anyons can only have multiplicity 0 or $1$ [101].
@@
We now apply the folding trick to obtain the SymTaco, as shown in Fig. 1, which results in a gapped boundary to vacuum for the $\mathcal{C}\boxtimes\overline{\mathcal{C}}$ TO. Importantly, anyons from this TO undergo a sequence of two condensations to get to the boundary: first, an anyon $a_{+}b_{-}\in\mathcal{C}\boxtimes\overline{\mathcal{C}}$ tunnels into the region with $\mathcal{C}^{\prime}\boxtimes\overline{\mathcal{C}}^{\prime}$ TO and turns into $\bigoplus_{\alpha\beta}W_{a\alpha}W_{b\beta}\,\alpha_{+}\beta_{-}$. Next, anyons from $\mathcal{C}^{\prime}\boxtimes\overline{\mathcal{C}}^{\prime}$ can condense at the gapped boundary to vacuum specified by condensation of the Lagrangian subalgebra
$\mathcal{L}^{\prime}=\bigoplus_{\alpha}\alpha_{+}\bar{\alpha}_{-}\,$, (40)
which corresponds to the canonical gapped boundary of the condensed TO $\mathcal{C}^{\prime}\boxtimes\overline{\mathcal{C}}^{\prime}$. In other words, the corresponding condensation matrix for this gapped boundary to vacuum is given by $M^{\prime}_{\alpha,\bar{\beta}}=\delta_{\alpha\beta}$. Hence, the gapped boundary to vacuum of the SymTaco, i.e., of the $\mathcal{C}\boxtimes\overline{\mathcal{C}}$ TO is specified by the data describing this sequence of condensations. The condensation matrix $M$ describing the folded boundary of the SymTaco  is therefore given by
$M_{a,\bar{b}}=W_{a,\alpha}W_{\bar{b},\bar{\beta}}\delta_{\alpha\bar{\beta}}=W_{a,\alpha}W_{\bar{b},\alpha}=(WW^{T})_{a\bar{b}}\,$, (41)
from which we see that $M_{a,\bar{b}}$ is a positive semidefinite matrix. Thus, the general condition for when a gapped boundary of an arbitrary TO $\mathcal{C}$ provides an admissible folded boundary of the corresponding SymTaco  is encoded in the positivity of the condensation matrix $M$, which is constructed through the intermediate step of the tunneling matrix $W$ (which is in turn specified by the condensable subalgebra $\mathcal{A}$ that is part of the data of the SymTaco  as defined in this paper).